OpenStudy (anonymous):

28.6% convert to lowest fraction

OpenStudy (mathmagician):

28.6%=286/1000=143/500

OpenStudy (anonymous):

% = /100 so 28.6% = 28.6/100 and multiply top and bottom by 10 to get 286/1000....